# Intelligent Polarforming Antenna Enhanced Sensing and Communication: Modeling and Optimization

### v1

@@ II-B IPA Channel Model
By combining (6) and (8), the steering vector of the $b$-th IPA subarray is given by
@@
where $\lambda$ denotes the carrier wavelength. Next, to determine effective antenna gain $g_{k,b}(\mathbf{u}_{b})$, we project $\mathbf{f}_{k}$ onto the local CCS of the $b$-th IPA subarray, denoted by
@@
Then, we represent $\tilde{\mathbf{f}}_{k,b}$ in the spherical coordinate system as $\tilde{\mathbf{f}}_{k,b}=[\cos(\tilde{\theta}_{k,b})\cos(\tilde{\phi}_{k,b}),\cos(\tilde{\theta}_{k,b})\sin(\tilde{\phi}_{k,b}),\sin(\tilde{\theta}_{k,b})]^{T}$, where $\tilde{\theta}_{k,b}$ and $\tilde{\phi}_{k,b}$ represent the corresponding directions of arrival (DoAs) in the local CCS (see Fig. 3). Finally, effective antenna gain $g_{k,b}(\mathbf{u}_{b})$ of the $b$-th IPA subarray along direction $(\tilde{\theta}_{k,b},\tilde{\phi}_{k,b})$ in the linear scale is given by
@@
which denotes the unpolarformed LoS channel between the $b$-th IPA subarray at BS and the $k$-th user. The term “unpolarformed” here refers to the fact that the polarforming effect is not considered yet. In (12), $\nu_{k}=\epsilon_{0}d_{k}^{-2}$ is the free-space path loss, where $\epsilon_{0}$ represents the channel power at the reference distance $d_{0}=1$ meter (m) [23], and $d_{k}>d_{0}$ denotes the distance between the $k$-th user’s location and the BS center.
@@
where $\mathbf{R}(\mathbf{u}_{b})\mathbf{e}_{\mathrm{v}}$ and $\mathbf{R}(\mathbf{u}_{b})\mathbf{e}_{\mathrm{h}}$ represent the mapping of the local CCS of the $\mathcal{V}$$\mathcal{H}$-element of the IPA subarray at the BS to the global CCS. Similarly, the receive field components are obtained by projecting the LoS signal direction onto the receive antenna using the projection matrix
@@
where $\mathbf{R}(\mathbf{u}_{k}^{\mathrm{r}})\mathbf{e}_{\mathrm{v}}$ and $\mathbf{R}(\mathbf{u}_{k}^{\mathrm{r}})\mathbf{e}_{\mathrm{h}}$ denote the transformation of the local CCS of the $\mathcal{V}$$\mathcal{H}$-element of the user’s IPA to the global CCS. Consequently, the dual-polarized response matrix between the $k$-th user and the dual-polarized antennas on the $b$-th IPA subarray at the BS is given by
@@
When the transmission distance is sufficiently large, the phase variation due to the LoS channel is the same for both $\mathcal{V}$- and $\mathcal{H}$-ports of the antenna, regardless of the pair of transmit and receive antennas. Therefore, the unpolarformed channel, $\overline{\mathbf{h}}_{k,b}(\mathbf{q}_{b},\mathbf{u}_{b})$, from the $\mathcal{V}$- and $\mathcal{H}$-ports of user $k$ to those of the $b$-th IPA subarray at the BS can be expressed as follows, using unpolarformed LoS channel $\mathbf{h}_{k,b}^{\mathrm{LoS}}(\mathbf{q}_{b},\mathbf{u}_{b})$ and dual-polarized response matrix $\mathbf{A}_{k,b}(\mathbf{u}_{b},\mathbf{u}_{k}^{\mathrm{r}})$:
@@
which captures the effects of wireless signal propagation across all potential polarization states. Different from the dual-polarized antenna [6], which requires two radio frequency (RF) chains since each antenna has two ports, each IPA is connected to a single RF chain for subsequent signal processing, thereby reducing power consumption (see Fig. 1). The $N$- dimensional IPA polarformed channel is then obtained by incorporating the polarization effect for each antenna through receive polarforming vector $\mathbf{w}_{k}$ and transmit polarforming vector $\mathbf{v}_{b}$, which is expressed as
@@
where $\mathbf{I}_{N}\otimes\mathbf{v}_{b}^{H}$ in (19) is due to the fact that all $N$ antennas in the same IPA subarray at BS use the same transmit polarforming vector $\mathbf{v}_{b}$. The polarforming vectors effectively combine the entries of channel $\overline{\mathbf{h}}_{k,b}(\mathbf{q}_{b},\mathbf{u}_{b})$ according to the polarization of each antenna at the transmitter/receiver.
Remark 1: The channel decomposition given in (20), introduced in this study for the first time, offers the advantage of decoupling the stable unpolarformed LoS channel component from the dynamic polarformed channel component due to transmitter/receiver polarforming and user rotations in 3D space. In this way, the unpolarformed channel remains unaffected when varying the polarforming vectors and user rotations. This distinctive IPA-specific channel representation can be utilized to design efficient localization algorithms that leverage the common parameters in the stable unpolarformed channel across different controllable polarforming vectors (see Section IV). Furthermore, the elegant structure of (20) facilitates the optimization of polarforming at both the transmitter and receiver that will be elaborated in Section V.
III Two-Timescale Protocol Design for IPA-Enhanced ISAC
Figure 4: Illustration of the proposed protocol for IPA-enhanced ISAC system.
In this section, we propose a two-timescale transmission protocol for the IPA-enhanced ISAC system, as illustrated in Fig. 4. Based on our proposed channel setup with users’ quasi-static locations and fast-varying rotations, we divide the user transmissions into multiple user location coherence time intervals, during each of which users’ locations are assumed to be constant. Each location coherence interval is further divided into two phases, namely, the slow-timescale phase for user localization and BS antenna position/rotation tuning, and the fast-timescale phase for polarforming adapting to users’ instantaneous orientations and their induced channel variations. Specifically, the detailed operation and design in these two phases are provided as follows.
  • Phase I (Slow Timescale Design) : During this initial phase, user locations are first sensed, and then BS antennas’ positions and rotations are determined based on the sensed user locations. To this end, the BS receives the pilot signals from the user devices which are transmitted using time-varying user polarforming vectors $\mathbf{w}_{k}$ to enable polarforming-based user localization at the BS (see Section IV). Then, based on the sensed user locations, we optimize the positions and rotations of all IPAs at the BS for maximizing the users’ average achievable rates (see Section V). The optimized antenna positions and rotations will then be implemented at the BS and remain unchanged during Phase II of each location coherence interval.
@@ III Two-Timescale Protocol Design for IPA-Enhanced ISAC
Note that by adapting to user rapid rotation with polarforming after antenna positions and rotations are set to their optimized values based on user locations, our scheme preserves the performance benefits of adapting to both the random and constant components of the IPA channel, thus achieving a good balance between performance and implementation cost.
@@ IV Polarforming-based Localization
In our proposed IPA-enhanced ISAC system, the accuracy of user localization significantly impacts the performance of polarforming and antenna position/rotation optimization. Once user location is determined, the stable unpolarformed channel component given in (20) between the user and all potential IPA positions and rotations at the BS can be reconstructed for slow-timescale antenna position and rotation optimization. Additionally, the user location-related dynamic polarformed channel component in (20) can be more efficiently acquired for fast-timescale polarforming optimization. Acquiring the stable unpolarformed channel and the dynamic polarformed channel components separately is a practically appealing approach for user localization, as these separate channels provide different information about user locations. Therefore, this section presents a polarforming-based user localization algorithm, where the user polarforming vectors dynamically vary according to a designed sequence to help extract the stable unpolarformed LoS channel components. The user locations are then derived based on the decoupled LoS channel components.
@@ IV-A Extraction of Unpolarformed Channel Component
To enable user localization sensing with a small number of BS antenna position-rotation placement pairs, IPA subarrays at the BS move across a set of $M$ different position-rotation pairs to collect pilot signals from all users. It is generally not feasible to sense the locations of all users widely distributed in the network using a single IPA subarray placed at a fixed position-rotation pair at the BS. This is due to the directional sparsity of IPA channels [22], where each user exhibits significant channel gains only with a small subset of IPA position-rotation pairs at the BS.
@@
with $\eta_{m,k,p}=\mathbf{v}_{m}^{H}\mathbf{A}_{m,k}(\mathbf{u}_{m},\mathbf{u}_{k}^{\mathrm{r}})\mathbf{w}_{k,p}$ as given in (20), for user $k$ at the $m$-th training position-rotation pair for IPA subarrays during pilot block $p$, with $p\in\{1,2,\dots,P\}$. In addition, $\mathbf{X}=[\mathbf{x}_{1},\cdots,\mathbf{x}_{K}]\in\mathbb{C}^{L\times K}$ denotes the horizontal stack of all pilots from all users, $\mathbf{H}_{m}=[\mathbf{h}_{m,1}^{\mathrm{LoS}}(\mathbf{q}_{m},\mathbf{u}_{m}),\mathbf{h}_{m,2}^{\mathrm{LoS}}(\mathbf{q}_{m},\mathbf{u}_{m}),\cdots,\mathbf{h}_{m,K}^{\mathrm{LoS}}(\mathbf{q}_{m},\mathbf{u}_{m})]^{T}\in\mathbb{C}^{K\times N},m\in\mathcal{M}$, denotes the collective unpolarformed LoS channel from all $K$ users to all antennas at the $m$-th IPA training position-rotation pair, and $\mathbf{W}_{m,p}\in\mathbb{C}^{L\times N}$ is the complex-valued additive white Gaussian noise (AWGN) matrix with independent and identically distributed (i.i.d.) entries following the distribution of $\mathcal{CN}(0,\sigma^{2})$.
@@ V Polarforming, Position, and Rotation Optimization
where $x_{k}\sim\mathcal{CN}(0,1)$ is the data symbol intended for the $k$-th user, $\mathbf{c}_{k}\in\mathbb{C}^{NB\times 1}$ represents the transmit precoder for user $k$, ${\mathbf{q}}=[\mathbf{q}_{1}^{T},\cdots,\mathbf{q}_{B}^{T}]^{T}\in\mathbb{R}^{3B\times 1},$ and $\mathbf{u}=[\mathbf{u}_{1}^{T},\mathbf{u}_{2}^{T},\cdots,\mathbf{u}_{B}^{T}]^{T}\in\mathbb{R}^{3B\times 1}$ are position and rotation vectors for all IPA subarrays at BS, and $n_{k}\sim\mathcal{CN}(0,\sigma^{2})$ is the AWGN. Using Shannon’s formula, the achievable rate of the $k$-th user over unit radio spectrum bandwidth is given by
@@
where $\mathbf{h}_{k}(\mathbf{q},\mathbf{u},\mathbf{w}_{k},\mathbf{v})=\left[\mathbf{h}_{k,1}^{T},\mathbf{h}_{k,2}^{T},\cdots,\mathbf{h}_{k,B}^{T}\right]^{T}\in\mathbb{C}^{NB\times 1}$ denotes the overall IPA channel between user $k$ and all the IPA subarrays at the BS, $\mathbf{v}=[\mathbf{v}_{1}^{T},\cdots,\mathbf{v}_{B}^{T}]^{T}$, and $\mathbf{c}=[\mathbf{c}_{1}^{T},\cdots,\mathbf{c}_{K}^{T}]^{T}$.
@@
where $\varrho_{k}$ represents the rate weight of user $k$, $\zeta$ is the total transmit power of the BS, the expectation is taken over the random channel variations due to arbitrary user rotations, and constraints (43b) and (43c) ensure that the receive and transmit polarforming vectors satisfy the discrete amplitude and phase requirements. Constraint (43e) ensures that the center of each IPA subarray is positioned within the convex 3D site space of the BS. As discussed in [9], minimum distance $d_{\min}$ enforced by constraint (43f) prevents overlapping and coupling between IPA subarrays. Constraint (43g) mitigates mutual signal reflections among BS antennas.
@@ V-B Slow Timescale Optimization
The slow-timescale optimization of antenna position vector $\mathbf{q}$ and rotation vector $\mathbf{u}$ in problem (P1) is a stochastic optimization problem due to the expectation operation in the objective function. Since this objective function is intractable, we approximate it as a deterministic function. Specifically, we independently generate $\overline{L}$ sets of random channel samples for all users and use their achievable rates averaged over these channel samples as a close approximation of the expected rate in the objective function of (P1), provided that $\overline{L}$ is sufficiently large. Given the user locations estimated at the beginning of Phase I of our proposed protocol, the BS can obtain the users’ LoS channels for all possible IPA position-rotation pairs at the BS and accordingly generate the $\overline{L}$ random channel samples by independently and randomly varying rotations $\mathbf{u}_{k}^{\mathrm{r}}$ of all users.
Suppose that the $l$-th sample of the IPA time-variant channels is ^$\hat{\mathcal{H}}^{l}=\{\mathbf{h}_{1}^{l},\mathbf{h}_{2}^{l},\cdots,\mathbf{h}_{K}^{l}\}$. The average sum rate of the $k$-th user is approximated by $\frac{1}{\overline{L}}\sum_{l=1}^{\overline{L}}R_{k}\left(\mathbf{q},\mathbf{u},\mathbf{w}_{k},\mathbf{v},\mathbf{c},\hat{\mathcal{H}}^{l}\right)$, where $\hat{\mathcal{H}}$ represents the collection of all $\overline{L}$ channel samples. Problem (P1) is thus recast as
@@
where $\tau$ is a positive penalty parameter, and $\mathcal{Q}(\mathbf{s})$ represents the penalty term for infeasible particles. Specifically, each element in $\mathcal{Q}(\mathbf{s})$ corresponds to position-rotation pair $(\mathbf{q},\mathbf{u})$ of IPA subarrays that violates constraints (43f) or (43g). Thus, $\mathcal{Q}(\mathbf{s})$ is defined as

### v2

@@ II-B PA Channel Model
By combining (6) and (8), the steering vector of the $b$-th PA subarray is given by
@@
where $\lambda$ denotes the carrier wavelength. Next, to determine effective antenna gain $g_{k,b}(\mathbf{u}_{b})$, we project $\mathbf{f}_{k}$ onto the local CCS of the $b$-th PA subarray, denoted by
@@
Then, we represent $\tilde{\mathbf{f}}_{k,b}$ in the spherical coordinate system as $\tilde{\mathbf{f}}_{k,b}=[\cos(\tilde{\theta}_{k,b})\cos(\tilde{\phi}_{k,b}),\cos(\tilde{\theta}_{k,b})\sin(\tilde{\phi}_{k,b}),\sin(\tilde{\theta}_{k,b})]^{T}$, where $\tilde{\theta}_{k,b}$ and $\tilde{\phi}_{k,b}$ represent the corresponding directions of arrival (DoAs) in the local CCS (see Fig. 3). Finally, effective antenna gain $g_{k,b}(\mathbf{u}_{b})$ of the $b$-th PA subarray along direction $(\tilde{\theta}_{k,b},\tilde{\phi}_{k,b})$ in the linear scale is given by
@@
which denotes the unpolarformed LoS channel between the $b$-th PA subarray at BS and the $k$-th user. The term “unpolarformed” here refers to the fact that the polarforming effect is not considered yet. In (12), $\nu_{k}=\epsilon_{0}d_{k}^{-2}$ is the free-space path loss, where $\epsilon_{0}$ represents the channel power at the reference distance $d_{0}=1$ meter (m) [24], and $d_{k}>d_{0}$ denotes the distance between the $k$-th user’s location and the BS center.
@@
where $\mathbf{R}(\mathbf{u}_{b})\mathbf{e}_{\mathrm{v}}$ and $\mathbf{R}(\mathbf{u}_{b})\mathbf{e}_{\mathrm{h}}$ represent the mapping of the local CCS of the $\mathcal{V}$$\mathcal{H}$-element of the PA subarray at the BS to the global CCS. Similarly, the receive field components are obtained by projecting the LoS signal direction onto the receive antenna using the projection matrix
@@
where $\mathbf{R}(\mathbf{u}_{k}^{\mathrm{r}})\mathbf{e}_{\mathrm{v}}$ and $\mathbf{R}(\mathbf{u}_{k}^{\mathrm{r}})\mathbf{e}_{\mathrm{h}}$ denote the transformation of the local CCS of the $\mathcal{V}$$\mathcal{H}$-element of the user’s PA to the global CCS. Consequently, the dual-polarized response matrix between the $k$-th user and the dual-polarized antennas on the $b$-th PA subarray at the BS is given by
@@
When the transmission distance is sufficiently large, the phase variation due to the LoS channel is the same for both $\mathcal{V}$- and $\mathcal{H}$-ports of the antenna, regardless of the pair of transmit and receive antennas. Therefore, the unpolarformed channel, $\overline{\mathbf{h}}_{k,b}(\mathbf{q}_{b},\mathbf{u}_{b})$, from the $\mathcal{V}$- and $\mathcal{H}$-ports of user $k$ to those of the $b$-th PA subarray at the BS can be expressed as follows, using unpolarformed LoS channel $\mathbf{h}_{k,b}^{\mathrm{LoS}}(\mathbf{q}_{b},\mathbf{u}_{b})$ and dual-polarized response matrix $\mathbf{A}_{k,b}(\mathbf{u}_{b},\mathbf{u}_{k}^{\mathrm{r}})$:
@@
which captures the effects of wireless signal propagation across all potential polarization states. Different from the dual-polarized antenna [6], which requires two radio frequency (RF) chains since each antenna has two ports, each PA is connected to a single RF chain for subsequent signal processing, thereby reducing power consumption (see Fig. 1). The $N$- dimensional PA polarformed channel is then obtained by incorporating the polarization effect for each antenna through receive polarforming vector $\mathbf{w}_{k}$ and transmit polarforming vector $\mathbf{v}_{b}$, which is expressed as
@@
where $\mathbf{I}_{N}\otimes\mathbf{v}_{b}^{H}$ in (19) is due to the fact that all $N$ antennas in the same PA subarray at BS use the same transmit polarforming vector $\mathbf{v}_{b}$. The polarforming vectors effectively combine the entries of channel $\overline{\mathbf{h}}_{k,b}(\mathbf{q}_{b},\mathbf{u}_{b})$ according to the polarization of each antenna at the transmitter/receiver.
Remark 1: The channel decomposition given in (20), introduced in this study for the first time, offers the advantage of decoupling the stable unpolarformed LoS channel component from the dynamic polarformed channel component due to transmitter/receiver polarforming and user rotations in 3D space. In this way, the unpolarformed channel remains unaffected when varying the polarforming vectors and user rotations. This distinctive PA-specific channel representation can be utilized to design efficient localization algorithms that leverage the common parameters in the stable unpolarformed channel across different controllable polarforming vectors (see Section IV). Furthermore, the elegant structure of (20) facilitates the optimization of polarforming at both the transmitter and receiver that will be elaborated in Section V.
III Two-Timescale Protocol Design for PA-Enhanced ISAC
Figure 4: Illustration of the proposed protocol for PA-enhanced ISAC system.
In this section, we propose a two-timescale transmission protocol for the PA-enhanced ISAC system, as illustrated in Fig. 4. Based on our proposed channel setup with users’ quasi-static locations and fast-varying rotations, we divide the user transmissions into multiple user location coherence time intervals, during each of which users’ locations are assumed to be constant. Each location coherence interval is further divided into two phases, namely, the slow-timescale phase for user localization and BS antenna position/rotation tuning, and the fast-timescale phase for polarforming adapting to users’ instantaneous orientations and their induced channel variations. Specifically, the detailed operation and design in these two phases are provided as follows.
  • Phase I (Slow Timescale Design) : During this initial phase, user locations are first sensed, and then BS antennas’ positions and rotations are determined based on the sensed user locations. To this end, the BS receives the pilot signals from the user devices which are transmitted using time-varying user polarforming vectors $\mathbf{w}_{k}$ to enable polarforming-based user localization at the BS (see Section IV). Then, based on the sensed user locations, we optimize the positions and rotations of all PAs at the BS for maximizing the users’ average achievable rates (see Section V). The optimized antenna positions and rotations will then be implemented at the BS and remain unchanged during Phase II of each location coherence interval.
@@ III Two-Timescale Protocol Design for PA-Enhanced ISAC
Note that by adapting to user rapid rotation with polarforming after antenna positions and rotations are set to their optimized values based on user locations, our scheme preserves the performance benefits of adapting to both the random and constant components of the PA channel, thus achieving a good balance between performance and implementation cost.
@@ IV Polarforming-based Localization
In our proposed PA-enhanced ISAC system, the accuracy of user localization significantly impacts the performance of polarforming and antenna position/rotation optimization. Once user location is determined, the stable unpolarformed channel component given in (20) between the user and all potential PA positions and rotations at the BS can be reconstructed for slow-timescale antenna position and rotation optimization. Additionally, the user location-related dynamic polarformed channel component in (20) can be more efficiently acquired for fast-timescale polarforming optimization. Acquiring the stable unpolarformed channel and the dynamic polarformed channel components separately is a practically appealing approach for user localization, as these separate channels provide different information about user locations. Therefore, this section presents a polarforming-based user localization algorithm, where the user polarforming vectors dynamically vary according to a designed sequence to help extract the stable unpolarformed LoS channel components. The user locations are then derived based on the decoupled LoS channel components.
@@ IV-A Extraction of Unpolarformed Channel Component
To enable user localization sensing with a small number of BS antenna position-rotation placement pairs, PA subarrays at the BS move across a set of $M$ different position-rotation pairs to collect pilot signals from all users. It is generally not feasible to sense the locations of all users widely distributed in the network using a single PA subarray placed at a fixed position-rotation pair at the BS. This is due to the directional sparsity of PA channels [23], where each user exhibits significant channel gains only with a small subset of PA position-rotation pairs at the BS.
@@
with $\eta_{m,k,p}=\mathbf{v}_{m}^{H}\mathbf{A}_{m,k}(\mathbf{u}_{m},\mathbf{u}_{k}^{\mathrm{r}})\mathbf{w}_{k,p}$ as given in (20), for user $k$ at the $m$-th training position-rotation pair for PA subarrays during pilot block $p$, with $p\in\{1,2,\dots,P\}$. In addition, $\mathbf{X}=[\mathbf{x}_{1},\cdots,\mathbf{x}_{K}]\in\mathbb{C}^{L\times K}$ denotes the horizontal stack of all pilots from all users, $\mathbf{H}_{m}=[\mathbf{h}_{m,1}^{\mathrm{LoS}}(\mathbf{q}_{m},\mathbf{u}_{m}),\mathbf{h}_{m,2}^{\mathrm{LoS}}(\mathbf{q}_{m},\mathbf{u}_{m}),\cdots,\mathbf{h}_{m,K}^{\mathrm{LoS}}(\mathbf{q}_{m},\mathbf{u}_{m})]^{T}\in\mathbb{C}^{K\times N},m\in\mathcal{M}$, denotes the collective unpolarformed LoS channel from all $K$ users to all antennas at the $m$-th PA training position-rotation pair, and $\mathbf{W}_{m,p}\in\mathbb{C}^{L\times N}$ is the complex-valued additive white Gaussian noise (AWGN) matrix with independent and identically distributed (i.i.d.) entries following the distribution of $\mathcal{CN}(0,\sigma^{2})$.
@@ V Polarforming, Position, and Rotation Optimization
where $x_{k}\sim\mathcal{CN}(0,1)$ is the data symbol intended for the $k$-th user, $\mathbf{c}_{k}\in\mathbb{C}^{NB\times 1}$ represents the transmit precoder for user $k$, ${\mathbf{q}}=[\mathbf{q}_{1}^{T},\cdots,\mathbf{q}_{B}^{T}]^{T}\in\mathbb{R}^{3B\times 1},$ and $\mathbf{u}=[\mathbf{u}_{1}^{T},\mathbf{u}_{2}^{T},\cdots,\mathbf{u}_{B}^{T}]^{T}\in\mathbb{R}^{3B\times 1}$ are position and rotation vectors for all PA subarrays at BS, and $n_{k}\sim\mathcal{CN}(0,\sigma^{2})$ is the AWGN. Using Shannon’s formula, the achievable rate of the $k$-th user over unit radio spectrum bandwidth is given by
@@
where $\mathbf{h}_{k}(\mathbf{q},\mathbf{u},\mathbf{w}_{k},\mathbf{v})=\left[\mathbf{h}_{k,1}^{T},\mathbf{h}_{k,2}^{T},\cdots,\mathbf{h}_{k,B}^{T}\right]^{T}\in\mathbb{C}^{NB\times 1}$ denotes the overall PA channel between user $k$ and all the PA subarrays at the BS, $\mathbf{v}=[\mathbf{v}_{1}^{T},\cdots,\mathbf{v}_{B}^{T}]^{T}$, and $\mathbf{c}=[\mathbf{c}_{1}^{T},\cdots,\mathbf{c}_{K}^{T}]^{T}$.
@@
where $\varrho_{k}$ represents the rate weight of user $k$, $\zeta$ is the total transmit power of the BS, the expectation is taken over the random channel variations due to arbitrary user rotations, and constraints (43b) and (43c) ensure that the receive and transmit polarforming vectors satisfy the discrete amplitude and phase requirements. Constraint (43e) ensures that the center of each PA subarray is positioned within the convex 3D site space of the BS. As discussed in [9], minimum distance $d_{\min}$ enforced by constraint (43f) prevents overlapping and coupling between PA subarrays. Constraint (43g) mitigates mutual signal reflections among BS antennas.
@@ V-B Slow Timescale Optimization
The slow-timescale optimization of antenna position vector $\mathbf{q}$ and rotation vector $\mathbf{u}$ in problem (P1) is a stochastic optimization problem due to the expectation operation in the objective function. Since this objective function is intractable, we approximate it as a deterministic function. Specifically, we independently generate $\overline{L}$ sets of random channel samples for all users and use their achievable rates averaged over these channel samples as a close approximation of the expected rate in the objective function of (P1), provided that $\overline{L}$ is sufficiently large. Given the user locations estimated at the beginning of Phase I of our proposed protocol, the BS can obtain the users’ LoS channels for all possible PA position-rotation pairs at the BS and accordingly generate the $\overline{L}$ random channel samples by independently and randomly varying rotations $\mathbf{u}_{k}^{\mathrm{r}}$ of all users.
Suppose that the $l$-th sample of the PA time-variant channels is ^$\hat{\mathcal{H}}^{l}=\{\mathbf{h}_{1}^{l},\mathbf{h}_{2}^{l},\cdots,\mathbf{h}_{K}^{l}\}$. The average sum rate of the $k$-th user is approximated by $\frac{1}{\overline{L}}\sum_{l=1}^{\overline{L}}R_{k}\left(\mathbf{q},\mathbf{u},\mathbf{w}_{k},\mathbf{v},\mathbf{c},\hat{\mathcal{H}}^{l}\right)$, where $\hat{\mathcal{H}}$ represents the collection of all $\overline{L}$ channel samples. Problem (P1) is thus recast as
@@
where $\tau$ is a positive penalty parameter, and $\mathcal{Q}(\mathbf{s})$ represents the penalty term for infeasible particles. Specifically, each element in $\mathcal{Q}(\mathbf{s})$ corresponds to position-rotation pair $(\mathbf{q},\mathbf{u})$ of PA subarrays that violates constraints (43f) [33, 34] or (43g) [9]. Thus, $\mathcal{Q}(\mathbf{s})$ is defined as